Distance=40miles, Time = 5 Minutes, Rate = 480mph
Using the formula D=R*T work through if you were missing one piece of information (Distance, Time, Rate)

Answer by stanbon(75887)

Distance=40miles, Time = 5 Minutes, Rate = 480mph
Using the formula D=R*T work through if you were missing one piece of information
----
distance = (1/12)(480) miles = 40 miles
----
rate = 40/(1/12) mph = 40*(12/1) = 480 mph
----
time = 40/480 hrs. = 4/48 = 1/12 hr = 5 minutes
=======================
Cheers,
Stan H.
===============